OpenStudy (anonymous):

Express the complex number in trigonometric form. -2 + 2square root of three i

OpenStudy (freckles):

\[-2+2 \sqrt{3} i \\ 2(-1+\sqrt{3} i)\] so let's look at just \[-1+\sqrt{3}i\] for now  you need to find theta there